OpenStudy (anonymous):

the sum of twice a number and 2 less than the number is the same as the difference between -18 and the number. What is the number?

OpenStudy (anonymous):

2x + (x - 2) = -18 - x

OpenStudy (anonymous):

thanxs

OpenStudy (anonymous):

3x - 2 = -18 - x add x to both sides 4x - 2 = -18 add 2 to both sides 4x = -16 divide 4 both sides x = -4